OpenStudy (anonymous):

the volme of a sphere is 5000pi=4/3pi r3(cubed) what is the surface area of the sphere?

OpenStudy (anonymous):

So 5000pi is the area?

OpenStudy (anonymous):

sorry volume

OpenStudy (mathstudent55):

For a circle: \(V = \dfrac{4}{3}\pi r^3\) In your case, \(V = 5000 \pi\) Set the volume formula equal to \(5000\pi\) and sove for r. Then use the formula for the surface area of a sphere: \(A = 4 \pi r^2\)